OpenStudy (phebe):

MEDALSS+FAN SIMPLIFY!! XD b. 4√10-√10

OpenStudy (anonymous):

\[4\sqrt{10}-\sqrt{10} \implies 4\sqrt{10}-1\sqrt{10} = 3\sqrt{10}\]

OpenStudy (phebe):

thanx

OpenStudy (anonymous):

Yw